OpenStudy (anonymous):

Check my answers please? 1. Write the equation of a polynomial that has zeros at –3 and 2. Equation: f(x) = (x+3)(x-2) --------------------------------------- Write each polynomial function in standard form. Then classify it by degree and by the number of terms. 2. n = 4m2 – m + 7m4 n = 4m3 + 4m – 2; cubic trinomial n = 7m4 + 4m2 – m; quartic trinomial---Correct n = 4m4 + 8m2 – m ; quartic trinomial n = 3m3 + 2m – 5; cubic trinomial ----------------------------------------- 3. f(t) = 4t + 3t3 + 2t – 7 f(t) = 3t3 + 6t – 7; cubic trinomial-----Correct f(t) = 2t3 + 4t –

OpenStudy (anonymous):

3. f(t) = 4t + 3t3 + 2t – 7 f(t) = 3t3 + 6t – 7; cubic trinomial-----Correct f(t) = 2t3 + 4t – 1; cubic trinomial f(t) = 5t3 + 2t – 7; cubic trinomial f(t) = 7t3 + 3t – 4; cubic trinomial ------------------------------------------ 4. f(r) = 5r + 7 + 2r2 f(r) = 8r3 + 5r + 1; cubic trinomial f(r) = r2 + 5r + 7; quadratic trinomial f(r) = 2r3 + 5r + 7; cubic trinomial f(r) = 2r2 + 5r + 7; quadratic trinomial----Correct ------------------------------------------ For each function, determine the zeros and their multiplicity. 5. y = (x + 2)^2(x – 5)^4 –2, multiplicity 2; 5, multiplicity 4----Correct 2, multiplicity 2; 5, multiplicity 3 –2, multiplicity 4; 5, multiplicity 2 2, multiplicity 2; 7, multiplicity 4

OpenStudy (anonymous):

6. y = (3x + 2)^3(x – 5)^5 4, multiplicity 2; 5, multiplicity 2 3, multiplicity 4; 2, multiplicity 4 -2/3, multiplicity 3; 5, multiplicity 5----Correct 2, multiplicity 3; 5, multiplicity 5 ------------------------------------------- 7. y = x^2(x + 4)^3(x – 1) 2, multiplicity 2; –4, multiplicity 2; 1, multiplicity 1 0, multiplicity 2; –4, multiplicity 3; 1, multiplicity 1----Correct 3, multiplicity 4; –3, multiplicity 2; 2, multiplicity 3 0, multiplicity 2; –3, multiplicity 2; 2, multiplicity 3

OpenStudy (zehanz):

1. Correct 2. Correct 3. Correct 4. Correct 5. Correct 6. Correct 7. Correct You leave nothing to correct for me ;( Congratulations!

OpenStudy (anonymous):

Can you help me with the last three questions? :) @ZeHanz

OpenStudy (zehanz):

They will be all correct again, I guess...

OpenStudy (anonymous):

no, i really don't understand them ^^"

OpenStudy (anonymous):

(x^3 + 3x^2 – x – 3) ÷ (x – 1) Find P(–3) for P(x) = –2x^4 – 3x^3 – x + 4.

OpenStudy (zehanz):

In the attachment, I would do a synthetic division: 4 | 2 -3 -18 -8 8 20 8 ________________ + 2 5 2 0 SO it would be (x-4)(2x²+5x+2) The first prob can also be done with synthetic division. P(-3): just substitute x=-3: -2(-3)^4-3(-3)^3-(-3)+4=... (I'm going to sleep now - it's 11:35 pm in Holland - got to work early tomorrow)